OpenStudy (anonymous):

solve 6x+10-2x=4(x +2)

OpenStudy (amorfide):

first of all the left side can be simplified by adding the x terms 6x-2x+10=4x+10 4x+10=4(x+2) now on the right hand side you want to multiply our the brackets 4(x+2)=4x+8 4x+10=4x+8 lol what? did you write the question right?

OpenStudy (anonymous):

yeah i did

OpenStudy (anonymous):

a. x=6/5 b. x=12 c. this is an identity and its true for all real numbers d. there is no solution these are all the possible anwsers

OpenStudy (amorfide):

there is no solution as when you simplify 4x+10=4x+8 you get 0=-2

OpenStudy (anonymous):

oh ok thank you

OpenStudy (andriod09):

@amorfide it wouldn't be -2, 10-8 is 2.

OpenStudy (amorfide):

it doesnt matter.... i can multiply both sides by -1 and get 0=2 there is still no solutions

OpenStudy (andriod09):

IK. I'm messing with you.

OpenStudy (amorfide):

shhh confusing me xD

OpenStudy (andriod09):

XD